# TextCrafter: Accurately Rendering Multiple Texts in Complex Visual Scenes

### v3

@@ 4.1 Benchmark
4.2 Evaluation Metrics and Baselines
The quality of CVTG are evaluated from five metrics: Word Accuracy and Normalized Edit Distance (NED) [32] are employed to evaluate the accuracy of text rendering. CLIPScore [12], VQAScore [22], and Aesthetics [20] are utilized to assess the quality of the generated images. For detailed configurations of the metrics calculation, please refer to the supplementary materials.
We compare our method on CVTG-2K with several state-of-the-art Text-to-Image models including Stable Diffusion 3.5 Large [9], FLUX.1 dev. To highlight the difference between CVTG and multi-instance generation tasks, we also compare with methods like 3DIS [45]. To highlight the distinction between CVTG and multi-line text generation tasks, we conducted comparative analyses of methods capable of generating multi-line text, such as AnyText [32] and TextDiffuser2 [5]. The parameter $\lambda$ described in Section 3.2 uses 0.4, and $r=5$ from Section 3.3. Notably, TextCrafter requires no training or fine-tuning and is compatible with different DiT backbone networks. We conducted experiments using both FLUX.1 dev and Stable Diffusion 3.5 Large as backbone networks. All experiments can be completed on a single A6000 GPU. For specific data on memory utilization and computational runtime, please refer to the supplementary materials.
@@ 4.2 Evaluation Metrics and Baselines
4.3 Quantitative Results
As demonstrated in Table 2, TextCrafter surpasses competing methods in both text accuracy and image quality on CVTG-2K, with TextCrafter(FLUX) improving word accuracy by over 45% relative to FLUX. While Stable Diffusion 3.5 and FLUX perform adequately in simple scenarios, their efficacy diminishes significantly with increased textual complexity. AnyText and TextDiffuser-2, despite being trained on rule-based data, fail to generalize to multi-region tasks, with AnyText performing 55% below TextCrafter even with improved mask coverage. Similarly, 3DIS struggles with visual text generation despite its multi-instance capabilities. As a training-free approach, TextCrafter maintains superior aesthetic scores compared to specialized alternatives while balancing textual accuracy with visual coherence, demonstrating exceptional robustness across complexity levels. Comprehensive results are available in the supplementary material.
4.4 Qualitative Results
Figure 6 illustrates comparative visual results between TextCrafter and state-of-the-art models (SD3.5, FLUX, AnyText, TextDiffuser-2, and 3DIS) on CVTG-2K. While SD3.5 and FLUX generate visually appealing images, they exhibit deficiencies in text rendering as regional complexity increases. AnyText demonstrates suboptimal performance with multi-word text instances, TextDiffuser-2 compromises background fidelity (resulting in diminished CLIPScore), and 3DIS deteriorates textual information during layout-to-depth conversion. Conversely, TextCrafter achieves superior visual harmony while accurately rendering multiple text regions without attribute confusion or prompt deviation. Additional qualitative evaluations are presented in the supplementary material.
4.5 Ablation Study
Table 3 presents quantitative ablation results for individual and combined components.
@@ 4.5 Ablation Study
Disabling instance fusion minimally impacts overall metrics but is vital for correct visual text localization; its absence often leads to misplaced text (e.g., "Hot Dogs," Figure 7, first column). Enabling it also suppresses hallucinated text, ensuring accurate placement (Figure 7, second column).
@@
Enabling region insulation improves all metrics by decoupling complex visual text, thereby reducing CVTG complexity and maximizing pre-trained model utility. It achieves over 60% word accuracy independently and reduces inter-text interference, enhancing clarity (e.g., "Arrival," Figure 7, fourth column).
@@
Text focus offers the most significant metric improvement, especially for small visual text, by enhancing clarity and accuracy (e.g., "Next Stop" and "VIP Access," Figure 7, fifth and last columns). Alone, it yields 63.51% word accuracy, roughly 10% below optimal, indicating a need for instance fusion and region insulation for stability in complex scenarios.
@@
The proposed Manhattan distance-based layout optimization significantly influences final generation by creating non-overlapping bounding boxes aligned with pretrained models’ spatial preferences. Table 3 quantitatively demonstrates its superiority over alternatives like Random or Fixed Layouts.
5 Conclusion
This paper addresses complex visual text generation (CVTG) by proposing TextCrafter, a novel training-free framework built on the DiT framework. It employs step-by-step refinement from prompts to clauses and finally to specific visual text. Key innovations include Instance Fusion for text-spatial carrier linkage, Region Insulation to reduce complexity and text loss, and a Text Focus approach for accurate small-text rendering. A new CVTG-2K benchmark dataset is also introduced. Extensive experiments demonstrate TextCrafter’s significant advantages over existing methods.
@@ 5 Conclusion
TextCrafter has two main limitations: (1) degraded performance with conflicting requirements (e.g., large text in small areas), where balancing spatial constraints and clarity is challenging; and (2) as a training-free method, its text enhancement can introduce artifacts that affect overall image aesthetics, especially in visually dense scenes. Future work will aim to develop mechanisms for high-fidelity text rendering while maintaining the aesthetic integrity of the image.

### v6

@@ Introduction
MARIOEval.
To supplement our evaluation of TextCrafter’s applicability in single-text and simple-scene scenarios, we selected a representative subset from the MARIOEval benchmark (Chen et al. 2024a) for comprehensive testing and analysis.
Metric AnyText TextDiffuser-2 TextCrafter
Word Accuracy↑ 0.1301 0.5466 0.6972
NED↑ 0.4401 0.7849 0.8855
CLIPScore↑ 0.7352 0.8240 0.8181
VQAScore↑ 0.7037 0.8669 0.8607
Aesthetics↑ 3.39 3.49 4.39
Table 3: Quantitative results on MARIOEval subset (Chen et al. 2024a). Bold values denote the best performance, while underlined values indicate the second-best performance for each metric.
Steps Word Accuracy↑ NED↑
Step1(Instance Fusion) 0.4422 0.6495
Step2(Region Insulation) 0.6116 0.7750
Step3(Text Focus) 0.6351 0.8055
Step1+Step2 0.6121 0.7762
Step2+Step3 0.7328 0.8651
All 0.7370 0.8679
Layout Methods Word Accuracy↑ NED↑
Random 0.6900 0.8374
Fixed 0.6994 0.8475
Optimizer 0.7370 0.8679
Table 4: Ablation studies on CVTG-2K dataset
Evaluation Metrics and Baselines
The experiments employ five metrics: Word Accuracy and Normalized Edit Distance (NED) (Tuo et al. 2024) are employed to evaluate the accuracy of text rendering. CLIPScore (Hessel et al. 2021), VQAScore (Lin et al. 2024), and Aesthetics (LAION-AI 2022) are utilized to assess the quality of the generated images. For detailed configurations of the metrics calculation, please refer to the supplementary materials. We compare our method on CVTG-2K with state-of-the-art Text-to-Image models including Stable Diffusion 3.5 Large (Esser et al. 2024), FLUX.1 dev. To highlight the distinction between CVTG and multi-line text generation, we compare with methods capable of generating multi-line text, such as AnyText (Tuo et al. 2024) and TextDiffuser2 (Chen et al. 2025). To highlight the difference between CVTG and multi-instance generation, we also compare with methods like RAG-Diffusion (Chen et al. 2024b) and 3DIS (Zhou et al. 2025b). Given that the MARIOEval benchmark exclusively comprises single visual text prompts without multi-instance configurations, we excluded RAG-Diffusion and 3DIS from the MARIOEval evaluation. Accordingly, AnyText and TextDiffuser serve as representative visual text generation baselines for MARIOEval evaluation.
The parameter $\lambda=0.4$ described in Instance Fusion, and $r=5$ in Region Insulation. Each baseline is run once on each data sample. Notably, TextCrafter is compatible with different DiT backbone networks, and we have implemented versions with both FLUX.1 and Stable Diffusion 3. All experiments can be completed on a single A6000 GPU. For specific data on memory utilization and computational runtime, please refer to the supplementary materials.
@@
Quantitative Results
As demonstrated in Table 2, TextCrafter surpasses competing methods in both text accuracy and image quality on CVTG-2K, with TextCrafter (FLUX) improving word accuracy by over 45% relative to FLUX. While Stable Diffusion 3.5 and FLUX perform adequately in simple scenarios, their efficacy diminishes significantly with increased textual complexity. AnyText and TextDiffuser-2, despite being trained on rule-based data, fail to generalize to multi-region tasks. Similarly, RAG-Diffusion and 3DIS struggles with visual text generation despite its multi-instance capabilities. While TextCrafter was not originally designed for single visual text scenarios, it nevertheless achieves the highest text rendering accuracy and aesthetic quality on established benchmarks such as MARIOEval while maintaining superior instruction-following capability, as shown in Table 3.
As a training-free approach, TextCrafter maintains superior aesthetic scores while balancing textual accuracy with visual coherence, demonstrating exceptional robustness across complexity levels. Comprehensive results are available in the supplementary material.
Qualitative Results
Figure 6 illustrates comparative visual results between TextCrafter and state-of-the-art models on CVTG-2K. While SD3.5 and FLUX generate visually appealing images, they exhibit deficiencies in text rendering as regional complexity increases. AnyText demonstrates suboptimal performance with multi-word text instances, TextDiffuser-2 compromises background fidelity (resulting in diminished CLIPScore), and 3DIS deteriorates textual information during layout-to-depth conversion. While RAG-Diffusion excels at region-aware generation, it tends to neglect the rendering of visual text. Conversely, TextCrafter achieves superior visual harmony while accurately rendering multiple text regions without attribute confusion or prompt deviation. To further validate the effectiveness, we conducted a user study with experienced image generation researchers, with detailed results provided in the supplementary material.
Ablation Study
Table 4 presents quantitative ablation results for individual and combined components. Figure 7 shows the qualitative results of missing a step and full mode.
@@
Disabling Instance Fusion minimally impacts overall metrics but is vital for correct visual text localization; its absence often leads to misplaced text (e.g., “Hot Dogs” Figure 7, first column). Enabling it also suppresses hallucinated text(e.g., Figure 7, second column). Despite an apparent performance drop when applied alone (Word Accuracy from 0.4965 to 0.4422), the model with all three components enabled still reaches state-of-the-art, emphasizing the progressive and synergistic role of Instance Fusion within the complete pipeline.
@@
Enabling Region Insulation improves all metrics by decoupling complex visual text, thereby reducing CVTG complexity and maximizing pre-trained model utility. It achieves over 60% word accuracy independently and reduces inter-text interference, enhancing clarity (e.g., “Arrival”, Figure 7, fourth column).
@@
Text Focus offers the most significant metric improvement, especially for small visual text(e.g., “Next Stop” and “VIP Access”, Figure 7, fifth and last columns). Alone, it yields 63.51% word accuracy, roughly 10% below optimal, indicating a need for Instance Fusion and Region Insulation for stability in complex scenarios.
@@
The proposed Manhattan distance-based layout optimizer significantly influences final generation by creating non-overlapping bounding boxes aligned with pretrained models’ spatial preferences. Table 4 quantitatively demonstrates its superiority over alternatives like Random or Fixed Layouts.
Conclusion
This paper addresses complex visual text generation (CVTG) by proposing TextCrafter, a novel training-free framework built on the DiT model. It employs step-by-step refinement from prompts to clauses and finally to specific visual text. Key innovations include Instance Fusion for text-spatial carrier linkage, Region Insulation to reduce complexity and text omission, and Text Focus for accurate small-text rendering. A new CVTG-2K benchmark dataset is also introduced. Extensive experiments demonstrate TextCrafter’s significant advantages over existing methods.
@@ Conclusion
TextCrafter is constrained by the generative capabilities of the base models and does not support visual text generation in languages other than English. We plan to address this limitation in future research.